What is the formula to determine a percentage increase of a given number.
For example: an insurance premium is $549.37, it renews at $595.21, by what percentage did it increase?

Answer by checkley77(12844)

1-(increased value/original value)Corrected. Sorry.
1-585.21/549.37=1-1.0834=.0834 or 8.34%.